the base of a triangle is 6 cm ,the opposite angle is 100 degree ,what is the max possible area??
Answer by Alan3354(69443)

the base of a triangle is 6 cm ,the opposite angle is 100 degree ,what is the max possible area??
---------------
Max area is an isosceles triangle --> the other 2 angles are 40 degrees.
h/3 = tan(40)
h = 3 cm * tan(40)
Area = b*h/2 = 9*tan(40)
=~ 7.552 sq units